# Matryoshka Quantization

### v1

@@ 5.2 Co-distillation for ${\rm MatQuant}$
Given the nested nature of the models trained using ${\rm MatQuant}$, we explored co-distillation, where the outputs from a higher-precision model are used as the target for the lower-precision nested model, either in a standalone fashion or alongside the ground truth target (weighted equally). Table 4 shows the effects of co-distillation applied to ${\rm MatQuant}$ with both OmniQuant and QAT on Gemma-2 9B. While int8 and int4 show no significant improvement, the nested int2 model benefits substantially from the int8 supervision, reaching $1.65\%$ higher accuracy than the non-co-distilled ${\rm MatQuant}$ with OmniQuant. This helps us push the int2 quantized Gemma-2 9B beyond $70\%$ average downstream accuracy for the first time across all our experiments. Co-distillation in ${\rm MatQuant}$ opens up avenues for interesting design choices that can further leverage the inherent nested structure of integer data types.
[…]
OmniQuant QAT
Data type Config. Task Avg. log pplx. Task Avg. log pplx.
int8 $[8,4,2]$ $\bf 73.97$ $\bf 2.451$ $74.77$ $2.301$
$[8,4,8\rightarrow 2]$ $73.40$ $2.467$ $74.72$ $\bf 2.298$
$[8,4,2,8\rightarrow 2]$ $73.46$ $2.466$ $74.62$ $2.299$
$[8,4,2,8\rightarrow 4;2]$ $73.32$ $2.466$ $\bf 74.80$ $2.302$
int4 $[8,4,2]$ $\bf 73.88$ $\bf 2.481$ $73.71$ $2.332$
$[8,4,8\rightarrow 2]$ $73.84$ $2.488$ $73.76$ $\bf 2.328$
$[8,4,2,8\rightarrow 2]$ $73.01$ $2.495$ $\bf 73.78$ $2.329$
$[8,4,2,8\rightarrow 4;2]$ $73.12$ $2.518$ $73.48$ $2.330$
int2 $[8,4,2]$ $68.52$ $2.809$ $62.32$ $2.756$
$[8,4,8\rightarrow 2]$ $69.2$ $2.796$ $61.81$ $\bf 2.740$
$[8,4,2,8\rightarrow 2]$ $\bf 70.17$ $\bf 2.778$ $\bf 62.51$ $2.746$
$[8,4,2,8\rightarrow 4;2]$ $69.72$ $2.804$ $62.12$ $2.746$
Table 5: Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ significantly improves upon the baseline for int2 and, at times, outperforms ${\rm MatQuant}$. Crucially, int8 and int4 performances of Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ experience a significant accuracy decrease (Tables 21 & 22).
@@ 5.3 Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$
In Tables 1 and 2, ${\rm MatQuant}$ performs on par with the explicitly trained baselines for int4, int8, and the interpolated int3 and int6 precisions. However, the int2 models show a significant accuracy improvement. To investigate this, we conducted a simple ablation in ${\rm MatQuant}$ by removing the loss terms for int4 and int8 (i.e., $R=\{2\}$ in Equation 7 or setting $\lambda_{4}=\lambda_{8}=0$) and present the results in Table 5. We call this version of ${\rm MatQuant}$ as Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$. With Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, we observe a further boost of up to $1.67\%$, in the accuracy of int2 models at a $\sim$2accuracy drop in the corresponding int4 and int8 models – int2 is still nested within int8. This improvement likely stems from the six additional bits available during ${\rm MatQuant}$-style training to optimize the int2 representation.
In the case of Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, gradient descent is free to tune these six additional bits to improve the overall quality of the int2 model. In ${\rm MatQuant}$, since we have additional losses to preserve the performance of the int4 and int8, the int2 performance is slightly worse than Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$. However, since the int4 and int8 models are typically very close in accuracy to the bfloat16 model, ${\rm MatQuant}$ can shift some of the weights to improve the int2 model. As int4 and int8 models have substantially more quantized buckets than int2, we hypothesize that shifting some weights into adjacent buckets may not significantly affect their performance; however, it can significantly impact int2’s performance. In fact, in the weight distributions presented in Fig 1c, we observe that ${\rm MatQuant}$ results in a model where larger number of weights are assigned to the higher-valued buckets. Conclusively, ${\rm MatQuant}$ and Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ inherently seem to be a better way of doing low-bit quantization.
@@
S.P. ${\rm MatQuant}$ $\bf 47.78$ $\bf 3.705$ $34.69$ $7.564$
${\rm MatQuant}$ $47.17$ $3.837$ $\bf 43.33$ $\bf 3.806$
int6 Sliced int8 $74.56$ $2.358$ $73.71$ $2.094$
@@
We present results for FFN + Attention quantization for QAT in Table 6. For int8, int4 and the interpolated int6 model, ${\rm MatQuant}$ performs on par with the Baseline. However, we found int2 and int3 to be very unstable while quantizing both, the FFN and the Attention parameters. Most recent works that do QAT for both the blocks Chen et al. (2024); Liu et al. (2024a); Du et al. (2024) either do some form of warm starting for the quantized parameters, or have additional distillation and auxiliary loss functions. In the naive setup of minimizing the loss with respect to the ground truth, we find QAT to be very unstable at lower precisions. However, both ${\rm MatQuant}$ and Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ are very stable further highlighting the benefits brought by ${\rm MatQuant}$ style training.
@@ 5.4 Deployment Considerations
Current hardware accelerators have native support for serving int8 and int4 quantized models. Additionally, custom-implemented CUDA kernels can can support various low-precision bit-widths, like int2 and int3 (Chee et al., 2024; Frantar et al., 2022). ${\rm MatQuant}$ can generate a large number of models at inference time. Depending on the serving environment, we can choose between Mix’n’Match models and homogeneous sliced models. For example, suppose the serving environment has a memory constraint equivalent to an int3 model but lacks optimized support for int3, while supporting int2. In this case, a Mix’n’Match model performing comparably to the int3 model could be deployed. More generally, as depicted in Figure 2, ${\rm MatQuant}$ densely spans the memory-versus-accuracy curve and can be leveraged to obtain the most performant model for a specific serving constraint. ${\rm MatQuant}$ can enable further research on hardware software co-design to effectively support elastic bit-widths on-the-fly during inference time.
@@ 6 Conclusions
In this work, we presented ${\rm MatQuant}$, a novel multi-scale training technique that leverages the nested structure of integer data types to simultaneously optimize model weight quantization across multiple precisions (int8, int4, and int2) within a single model. This general-purpose method, applicable to learning-based quantization techniques like OmniQuant and QAT, produces models with comparable accuracy to baselines for int8 and int4, while achieving significant improvements, up to $10\%$ (using co-distillation), for int2 models. ${\rm MatQuant}$ further enables bit-width interpolation and layer-wise mix-and-match for flexible accuracy-cost trade-offs, promising more efficient deployment of large models across various hardware settings. Finally, ${\rm MatQuant}$ also helped discover Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, which significantly improves standalone low-bit quantization.
@@ Acknowledgments
We are grateful to Varun Yerram, Shreya Pathak and Devvrit for assistance in setting up inference pipelines, Praneeth Netrapalli, Rakesh Shivanna, Tom Duerig, Abhijit Ogale, Jon Shlens, Ali Farhadi and Rahul Sukthankar for helpful discussions, support and feedback.
Appendix A Addition Training Details
We run all our experiments on TPUv5e chips. For OmniQuant experiments, we use a constant learning rate of $31e-$ 3 and for QAT experiments, we linearly warmup the learning rate to $51e-$ 5 for 150 and use a consine decay schedule thereafter. For OmniQuant experiments, we sample 128 examples with a sequence length of 2048 from the C4 dataset (Raffel et al., 2020) and train using a batch size of 4. We train for a total of 10M tokens for all models except the int2 baseline, where we train the model for 20M tokens (Shao et al., 2023). For QAT experiments, we sample a fixed set of 100M tokens from the C4 dataset and train all our models using a batch size of 16 and a sequence length of 8192 for a single epoch. For Attn + FFN experiments with QAT, we sample a fixed set of 300M tokens from C4 and train with a batch size of 16 for a single epoch.

### v2

@@ 5.2 Co-distillation for ${\rm MatQuant}$
Given the nested nature of the models trained using ${\rm MatQuant}$, we explored co-distillation, where the outputs from a higher-precision model are used as the target for the lower-precision nested model, either in a standalone fashion or alongside the ground truth target (weighted equally). Table 4 shows the effects of co-distillation applied to ${\rm MatQuant}$ with both OmniQuant and QAT on Gemma-2 9B. While int8 and int4 show no significant improvement, the nested int2 model benefits substantially from the int8 supervision, reaching $0.97\%$ higher accuracy than the non-co-distilled ${\rm MatQuant}$ with OmniQuant. Co-distillation in ${\rm MatQuant}$ opens up avenues for interesting design choices that can further leverage the inherent nested structure of integer data types.
Table 5: Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ significantly improves upon the baseline for int2 and, at times, outperforms ${\rm MatQuant}$. Crucially, int8 and int4 performances of Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ experience a significant accuracy decrease (as shown in Tables 23 & 24) in Appendix G).
@@ 5.3 Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$
In Tables 1 and 2, ${\rm MatQuant}$ performs on par with the explicitly trained baselines for int4, int8, and the interpolated int3 and int6 precisions. However, the int2 models show a significant accuracy improvement. To investigate this, we conducted a simple ablation in ${\rm MatQuant}$ by removing the loss terms for int4 and int8 (i.e., $R=\{2\}$ in Equation 7 or setting $\lambda_{4}=\lambda_{8}=0$) and present the results in Table 5. We call this version of ${\rm MatQuant}$ as Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$. With Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, we observe a further boost of up to $1.05\%$, in the accuracy of int2 models at a $\sim$2accuracy drop in the corresponding int4 and int8 models – int2 is still nested within int8. This improvement likely stems from the six additional bits available during ${\rm MatQuant}$-style training to optimize the int2 representation.
In the case of Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, gradient descent is free to tune these six additional bits to improve the overall quality of the int2 model. In ${\rm MatQuant}$, since we have additional losses to preserve the performance of the int4 and int8, the int2 performance is slightly worse than Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$. However, since the int4 and int8 models are typically very close in accuracy to the bfloat16 model, ${\rm MatQuant}$ can shift some of the weights to improve the int2 model. As int4 and int8 models have substantially more quantized buckets than int2, we hypothesize that shifting some weights into adjacent buckets may not significantly affect their performance; however, it can significantly impact int2’s performance. In fact, in the weight distributions presented in Fig 1c, we observe that ${\rm MatQuant}$ results in a model where larger number of weights are assigned to the higher-valued buckets. Conclusively, ${\rm MatQuant}$ and Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ inherently seem to be a better way of performing low-bit quantization.
@@
S.P. ${\rm MatQuant}$ $\bf 45.69$ $\bf 3.780$ $35.35$ $7.761$
${\rm MatQuant}$ $44.19$ $3.826$ $\bf 38.36$ $\bf 10.971$
int6 Sliced int8 $74.49$ $2.290$ $73.61$ $2.104$
@@
We present results for FFN + Attention quantization for QAT in Table 6. For int8, int4 and the interpolated int6 model, ${\rm MatQuant}$ performs on par with the Baseline. However, we found int2 and int3 to be very unstable while quantizing both, the FFN and the Attention parameters. Most recent works that do QAT for both the blocks Chen et al. (2024); Liu et al. (2024a); Du et al. (2024) either do some form of warm starting for the quantized parameters, or have additional distillation and auxiliary loss functions. In the naive setup of minimizing the loss with respect to the ground truth, we find QAT to be very unstable at lower precisions. On the other hand, both ${\rm MatQuant}$ and Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$ are very stable further highlighting the benefits brought by ${\rm MatQuant}$ style training.
@@ 5.4 Deployment Considerations
Current hardware accelerators have native support for serving int8 and int4 quantized models. Additionally, custom-implemented CUDA kernels can can support various low-precision bit-widths, like int2 and int3 (Chee et al., 2024; Frantar et al., 2022). ${\rm MatQuant}$ can generate a large number of models at inference time. Depending on the serving environment, we can choose between Mix’n’Match models and homogeneous sliced models. For example, suppose the serving environment has a memory constraint equivalent to an int3 model but lacks optimized support for int3, while supporting int2. In this case, a Mix’n’Match model with a small performance drop when compared to the sliced int3 model could be deployed. More generally, as depicted in Figure 2, ${\rm MatQuant}$ densely spans the memory-versus-accuracy curve and can be leveraged to obtain performant model for several serving constraints. ${\rm MatQuant}$ can enable further research on hardware software co-design to effectively support elastic bit-widths on-the-fly during inference.
@@ 5.5 Extension to Floating Point
Table 7: Results comparing ${\rm MatQuant}$ with Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ for Gemma-2 2B, 9B, and Mistral 7B, with OmniQuant as the base algorithm. We find that for the 2-bit model, having an extra bucket significantly boosts the performance, however, this is not the case with the higher precisions.
Method Gemma-2 2B Gemma-2 9B Mistral 7B
OmniQuant Avg. Bits Task Avg. log pplx. Avg. Bits Task Avg. log pplx. Avg. Bits Task Avg. log pplx.
bfloat16 $68.21$ $2.551$ $74.38$ $2.418$ $73.99$ $2.110$
${\rm MatQuant}$ $8$ $68.02$ $2.570$ $8$ $74.05$ $2.438$ $8$ $73.65$ $2.125$
Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $8$ $67.85$ $2.580$ $8$ $74.33$ $2.446$ $8$ $73.46$ $2.132$
${\rm MatQuant}$ $4$ $66.58$ $2.618$ $4$ $73.83$ $2.491$ $4$ $73.06$ $2.153$
Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $4.023$ $66.54$ $2.617$ $4.022$ $74.26$ $2.470$ $4.022$ $73.13$ $2.155$
${\rm MatQuant}$ $2$ $\bf 52.37$ $\bf 3.800$ $2$ $\bf 63.35$ $\bf 3.187$ $2$ $\bf 62.75$ $\bf 3.153$
Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $2.052$ $\bf 55.70$ $\bf 3.355$ $2.050$ $\bf 68.25$ $\bf 2.823$ $2.051$ $\bf 65.99$ $\bf 2.569$
${\rm MatQuant}$ $6$ $67.52$ $2.574$ $6$ $73.92$ $2.440$ $6$ $73.63$ $2.127$
Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $6.018$ $68.01$ $2.582$ $6.018$ $74.50$ $2.446$ $6.018$ $73.59$ $2.139$
${\rm MatQuant}$ $3$ $64.47$ $2.618$ $3$ $72.87$ $2.607$ $3$ $71.16$ $2.238$
Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ $3.031$ $63.24$ $2.757$ $3.029$ $73.25$ $2.535$ $3.030$ $71.55$ $2.228$
@@ 6 Conclusions
In this work, we presented ${\rm MatQuant}$, a novel multi-scale training technique that leverages the nested structure of integer data types to simultaneously optimize model weight quantization across multiple precisions (int8, int4, and int2) within a single model. This general-purpose method, applicable to learning-based quantization techniques like OmniQuant and QAT, produces models with comparable accuracy to baselines for int8 and int4, while achieving significant improvements, up to $7\%$ for int2 models. ${\rm MatQuant}$ further enables bit-width interpolation and layer-wise mix-and-match for flexible accuracy-cost trade-offs, promising more efficient deployment of large models across various hardware settings. Finally, ${\rm MatQuant}$ also helped discover Single Precison ${\rm Single\text{ }Precison\text{ }MatQuant}$, which significantly improves standalone low-bit quantization.
[…]
Gemma-2 9B OmniQuant
${\rm MatQuant}$ E.P. ${\rm MatQuant}$
Avg. Bits Config. Task Avg. log pplx. Task Avg. log pplx.
$\left(8,8\right)$ $[8,4,2]$ $74.05$ $2.438$ $73.97$ $2.451$
$[8,4,8\rightarrow 2]$ $72.76$ $2.473$ $73.40$ $2.467$
$[8,4,2,8\rightarrow 2]$ $73.99$ $2.435$ $73.46$ $2.466$
$[8,4,2,8\rightarrow 4;2]$ $73.85$ $2.437$ $73.32$ $2.466$
$\left(4,4.022\right)$ $[8,4,2]$ $73.83$ $2.491$ $73.88$ $2.481$
$[8,4,8\rightarrow 2]$ $72.65$ $2.519$ $73.84$ $2.488$
$[8,4,2,8\rightarrow 2]$ $73.63$ $2.486$ $73.01$ $2.495$
$[8,4,2,8\rightarrow 4;2]$ $73.55$ $2.478$ $73.12$ $2.518$
$\left(2,2.050\right)$ $[8,4,2]$ $63.35$ $\bf 3.187$ $68.52$ $2.809$
$[8,4,8\rightarrow 2]$ $62.64$ $3.289$ $69.2$ $2.796$
$[8,4,2,8\rightarrow 2]$ $62.91$ $3.138$ $\bf 70.17$ $\bf 2.778$
$[8,4,2,8\rightarrow 4;2]$ $\bf 64.32$ $3.227$ $69.72$ $2.804$
7 Errata
In the first draft of the paper, we had a bug and used the following equation to train and quantize our models:
Equation 8 clearly allows an extra bucket to be included into the quantization range, i.e, a $r$-bit model would have $12^{r}+$ 1 possible values instead of $2^{r}$. For example, consider slicing the first two MSBs from an unsigned int8 value, $234$. As per Equation 6, $234$ first gets rounded to $4$, following which it gets clipped to $3$, and finally is scaled up to $3*64=192$ (Note that ${\rm MatQuant}$ int2 allows for 00, $64$, $128$, $192$). However, since the clipping operation is missing in Equation 8, $4$ is never clipped down to $3$, and $S(q^{c},r)$ is now $4*64=256$ Thus, for certain int2 values in our final quantized model, we will have to store an extra bit. This is the case with int3, int4 and int6 as well where an extra bit is required to represent certain values. In Table 7, we can see that the fraction of parameters that fall into this extra bucket is very small. However, for our 2-bit models, this additional bucket gives significant improvements in performance, for example, in Table 7 int2 Gemma-2 9B’s average downstream accuracy goes up by $5$% when trained with an additional bucket (referred to as Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ in Table 7). This number is further boosted to $6$% with co-distillation, as evidenced by Table 8. We hypothesize that this additional bucket helps with capturing the outliers and thus leads to a significant performance boost. As highlighted by recent work (Dettmers et al., 2023; Kim et al., 2024), it is crucial to store certain outliers full precision. Interestingly, we show that even a single bit is enough to capture several of these outliers, especially for low bit quantization. Finally, note that this performance boost is not very evident in higher precisions where there are enough buckets to account for the outliers.
Figure 3: Mix’n’Match on Gemma-2 9B model trained using Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ with OmniQuant as the base algorithm allows elastic pareto-optimal accuracy-vs-cost model extraction for free during deployment.
Mix’n’Match
As shown in Figure 3 with a strong int2 model (i.e., 2.050 bits on average), Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ Mix’n’Match densely spans the Pareto-optimal accuracy-vs-bits-per-FFN-parameter (memory/cost) trade-off for Gemma-2 9B model trained using MatQuant with Omni-Quant – sometimes even improving on the bfloat16 model accuracy. Consequently, hardware supporting only int2 and int4 data types can still accommodate a model with a memory footprint similar to that of an int3 quantized model, and quality comparable or superior to int3; the additional bits required in the case of int2 can be packed into int2/int4. However, custom CUDA kernel would be required to enable sparse additions of these additional bits to the model weights.
Impact Statement
This paper introduces a novel technique designed to advance the field of machine learning, specifically in the domain of model compression and efficient deployment for large language models. By enabling the creation of versatile, multi-scale models that can operate across various bit-widths, our work has the potential to democratize access to these powerful technologies by making them more resource-efficient and deployable on a wider range of hardware. This could lead to positive impacts such as more sustainable AI systems and greater accessibility for users with limited computational resources. While there are potential risks associated with the broad deployment of powerful AI systems, these are not unique to our work, and we believe the benefits of efficient and accessible AI through innovations like ${\rm MatQuant}$ have significant potential for societal good. We encourage further investigation into how novel quantization techniques can play a role in future sustainable AI development.
@@ Acknowledgments
We are grateful to Varun Yerram, Shreya Pathak and Devvrit for assistance in setting up inference pipelines, Shivani Agrawal. Utku Evci, Praneeth Netrapalli, Rakesh Shivanna, Tom Duerig, Abhijit Ogale, Jon Shlens, Ali Farhadi and Rahul Sukthankar for helpful discussions, support and feedback.
Appendix A Particulars of the Slicing Operation.
To extract a $r$-bit model from a $c$-bit model, we start by slicing out the most significant $r-1$ bits. We use $1$ for the $r^{\text{th}}$ bit if the $(r+1)^{\text{th}}$, else, we use 00. This is captured by the round function in Equation 6 and is done to push values to higher buckets as we expect them to be more informative (Sun et al., 2024). For example, consider the the unsigned int8 value $53$. The first two MSBs are 00s. Naively slicing them would round down $53$ to 00, however, we want to round it up to $1$. Since the bit corresponding to $32$ is set, i.e., the $(r+1)^{\text{th}}$ MSB, instead of rounding $53$ down to 00, we round it up to $1$.
The $\text{clamp}(\cdot)$ operation is also equally important. The rounding operation in Equation 6 will round $240$ down to $4$, however, unsigned int2 operates with only $0,1,2,3$. $\text{clamp}(\cdot)$ here would make sure that $4$ is clamped down to $3$.
Appendix B Addition Training Details
We run all our experiments on TPUv5e chips. For OmniQuant experiments, we use a constant learning rate of $31e-$ 3 and for QAT experiments, we linearly warmup the learning rate to $51e-$ 5 for 150 and use a consine decay schedule thereafter. For OmniQuant experiments, we sample 128 examples with a sequence length of 2048 from the C4 dataset (Raffel et al., 2020) and train using a batch size of 4. We train for a total of 10M tokens for all models except the int2 baseline, where we train the model for 20M tokens (Shao et al., 2023). For Co-distillation experiments where OmniQuant is the base algorithm, we train for a total of 8.3M tokens. For QAT experiments, we sample a fixed set of 100M tokens from the C4 dataset and train all our models using a batch size of 16 and a sequence length of 8192 for a single epoch. For Attn + FFN experiments with QAT, we sample a fixed set of 300M tokens from C4 and train with a batch size of 16 for a single epoch. We use $(\lambda_{8},\lambda_{4},\lambda_{2})=(0.1,0.1,1.0)$ for all our Gemma experiments unless otherwise stated. In the case of Mistral 7B, for OmniQuant experiments, we use $(\lambda_{8},\lambda_{4},\lambda_{2})=(0.4,0.4,1.0)$, and for QAT experiments we use $(\lambda_{8},\lambda_{4},\lambda_{2})=(0.2,0.2,1.0)$. For all our Extra Precison ${\rm Extra\text{ }Precison\text{ }MatQuant}$ experiments, we use $(\lambda_{8},\lambda_{4},\lambda_{2})=(1.0,1.0,1.0)$.